# Gravothermalizing into primordial black holes, boson stars, and cannibal stars

### v2

@@ VI.1 Minimal gravothermal accretion
Within the entire parameter region where a collapse takes place, we find that cannibal annhilations cannot be ignored. This can be seen by comparing the condition for negligible cannibal annihilations in eq. (74) with eq. (III.2): The necessary condition for the two to be compatible can be obtained by using just the first term in eq. (III.2) as,
@@
The white-dashed countours show the PBH abundance relative to dark matter, $f_{\rm BH}$, in base 10 logarithm. Since $f_{\rm BH}$ scales as $M_{\rm BH}^{-2.5}$ (see eq. (101)) and thus the total abundance is dominated by the lightest PBHs, the $f_{\rm BH}$ contours are shown for the minimum mass $M_{-}$. The dark blue region shows where $f_{\mathrm{BH}}>1$; in this region one sees that the minimum mass $M_{-}$ is larger than $10^{15}\,\mathrm{g}$, hence all PBHs survive until today and overdominate the universe. On the other hand, in the region with $M_{-}\lesssim 10^{15}\,\mathrm{g}$, which appears for $\lambda=1$, at least part of the PBHs evaporate and are subject to further observational constraints. The region labeled “CMB+BBN” shows where the PBH spectrum intersects the constraints from CMB and/or BBN (see fig. 2) and is therefore excluded.
@@ VI.2 Maximal gravothermal accretion
With accretion, PBHs form for a wide range of EMDE parameters in both cases with and without cannibalism, and even with $\lambda$ as small as $10^{-5}$. The rather efficient PBH production, however, leads to PBH overabundance or exclusion by CMB/BBN limits in a sizeable parameter region. The upper right edge of the “CMB+BBN” exclusion region more or less overlaps with the $M_{+}=10^{9}\,\mathrm{g}$ contour. Thus, the produced PBHs are so abundant that almost all PBHs that evaporate after the onset of BBN are ruled out. In figs. 4(a)4(c), the lower left part of the “CMB+BBN” region overlaps with “$f_{\mathrm{BH}}>1$,” with the former displayed in front of the latter. Hence, there is no place where the PBHs can explain the entire dark matter while being consistent with observations (except for with fine-tuned parameters on the left edge of the “$f_{\mathrm{BH}}>1$” region).
@@ VIII Summary and conclusion
The formation of cannibal stars is a particularly salient feature of the gravothermal catastrophe since the self-interactions inducing the thermalization also enable the number-changing interactions which produce heat to support the stars. For the toy EMDE model with a $\phi^{4}$ interaction, we found that the gravothermal catastrophe predominantly forms cannibal stars. If the stars accrete the surrounding particles, they can eventually collapse into PBHs. Alternatively, in extended models where the number-changing interactions are forbidden, PBHs can form from the direct gravothermal collapse of halos.
@@ Acknowledgements.
We are thankful to Yiming Zhong, Wei-Xiang Feng, Petr Tinyakov, Tomas Reis, Aleksandr Azatov, Kiyotomo Ichiki, Donghui Jeong, Mak Pavičević, Serguey Petcov, and Emiliano Sefusatti for useful discussions. This research was supported in part by INFN TASP. T.K. also acknowledges support from the European Union - NextGenerationEU through the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W), and from JSPS KAKENHI (JP22K03595). D.P. was partially supported by the National Science Centre, Poland, under research grant no. 2020/38/E/ST2/00243.
@@ Appendix B Evolution of halos using Press–Schechter formalism
There are only two terms on the right-hand side with time dependence: $\bar{\rho}_{m}$ and $\sigma_{\rm R}$. The matter density simply evolves as $\bar{\rho}_{m}\propto 1/a^{3}$. Since during the matter-dominated era, the matter density perturbations grow proportional to the scale factor (cf. eq. (7)), we find that $\sigma_{\rm R}\propto a$. Thus, $dn_{\mathrm{h}}/dM\propto 1/a^{4}$, or equivalently, the comoving number density of halos falls as $1/a$.

### v3

@@ VI.1 Minimal gravothermal accretion
Within the entire parameter region where a collapse takes place, we find that cannibal annihilations cannot be ignored. This can be seen by comparing the condition for negligible cannibal annihilations in eq. (74) with eq. (III.2): The necessary condition for the two to be compatible can be obtained by using just the first term in eq. (III.2) as,
@@
The white-dashed contours show the PBH abundance relative to dark matter, $f_{\rm BH}$, in base 10 logarithm. Since $f_{\rm BH}$ scales as $M_{\rm BH}^{-2.5}$ (see eq. (101)) and thus the total abundance is dominated by the lightest PBHs, the $f_{\rm BH}$ contours are shown for the minimum mass $M_{-}$. The dark blue region shows where $f_{\mathrm{BH}}>1$; in this region one sees that the minimum mass $M_{-}$ is larger than $10^{15}\,\mathrm{g}$, hence all PBHs survive until today and overdominate the universe. On the other hand, in the region with $M_{-}\lesssim 10^{15}\,\mathrm{g}$, which appears for $\lambda=1$, at least part of the PBHs evaporate and are subject to further observational constraints. The region labeled “CMB+BBN” shows where the PBH spectrum intersects the constraints from CMB and/or BBN (see fig. 2) and is therefore excluded.
@@ VI.2 Maximal gravothermal accretion
With accretion, PBHs form for a wide range of EMDE parameters in both cases with and without cannibalism, and even with $\lambda$ as small as $10^{-5}$. The rather efficient PBH production, however, leads to PBH overabundance or exclusion by CMB/BBN limits in a sizeable parameter region. The upper right edge of the “CMB+BBN” exclusion region more or less overlaps with the $M_{+}=10^{9}\,\mathrm{g}$ contour. Thus, the produced PBHs are so abundant that almost all PBHs that evaporate after the onset of BBN are ruled out. In figs. 4(a)4(c), the lower left part of the “CMB+BBN” region overlaps with $f_{\mathrm{BH}}>1$,” with the former displayed in front of the latter. Hence, there is no place where the PBHs can explain the entire dark matter while being consistent with observations (except for with fine-tuned parameters on the left edge of the $f_{\mathrm{BH}}>1$” region).
@@ VIII Summary and conclusion
The formation of cannibal stars is a particularly salient feature of the gravothermal catastrophe since the self-interactions inducing the thermalization also enable the number-changing interactions that produce heat to support the stars. For the toy EMDE model with a $\phi^{4}$ interaction, we found that the gravothermal catastrophe predominantly forms cannibal stars. If the stars accrete the surrounding particles, they can eventually collapse into PBHs. Alternatively, in extended models where the number-changing interactions are forbidden, PBHs can form from the direct gravothermal collapse of halos.
@@ Acknowledgements.
We are thankful to Yiming Zhong, Wei-Xiang Feng, Petr Tinyakov, Tomas Reis, Aleksandr Azatov, Kiyotomo Ichiki, Donghui Jeong, Pierluigi Monaco, Aseem Paranjape, Mak Pavičević, Serguey Petcov, and Emiliano Sefusatti for useful discussions. This research was supported in part by INFN TASP. T.K. also acknowledges support from the European Union - NextGenerationEU through the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W), and from JSPS KAKENHI (JP22K03595). D.P. was partially supported by the National Science Centre, Poland, under research grant no. 2020/38/E/ST2/00243.
@@ Appendix B Evolution of halos using Press–Schechter formalism
There are only two terms on the right-hand side with time dependence: $\bar{\rho}_{m}$ and $\sigma_{\rm R}$. The matter density simply evolves as $\bar{\rho}_{m}\propto 1/a^{3}$. Since during the matter-dominated era, the matter density perturbations grow proportional to the scale factor (cf. eq. (7)), we find that $\sigma_{\rm R}\propto a$. Thus, $dn_{\mathrm{h}}/dM\propto 1/a^{4}$, or equivalently, the comoving number density of halos [^27] Crudely, one considers the increase in time of $dn_{\mathrm{h}}/dM$ due to the factor $\exp(-\delta_{c}^{2}/2\sigma_{\mathrm{R}}^{2})$ to correspond to the formation of halos, and the decrease due to $\delta_{c}/\sigma_{\mathrm{R}}$ as accretion into larger halos (here $d\ln\sigma_{\mathrm{R}}/dM$ is considered to be constant). This suggests that most of the halo formation takes place during the linear regime, i.e., $\sigma_{\mathrm{R}}\lesssim\delta_{c}$. falls as $1/a$.